# J-HERTz: J-PLUS Heritage Exploration of Radio Targets at z 5

### v1

@@ 1 Introduction
Radio astronomy has made substantial advances since the first radio emission coming from the center of our galaxy was discovered (Jansky1933). In 90 years, radio continuum surveys have improved from angular resolutions of several degrees to sub-arcsecond scales thanks to the development of radio interferometry, and from frequencies of a few hundred MHz to a full coverage of the radio window up to millimeter wavelengths. This advancement came hand in hand with the technical development that enabled the construction of larger and more efficient antennas, as well as improved backend systems with more sensitive receivers and faster correlators. Some of the most important radio continuum surveys in the last decades include the NVSS111NRAO (National Radio Astronomy Observatory) VLA Sky Survey. (Condon1998) at 1.4 GHz, the Cambridge 3C Radio Survey (Bennett1962) at 178 MHz, the Very Large Array (VLA) FIRST222Faint Images of the Radio Sky at Twenty-Centimeters. Survey (Becker1994) at 1.5 GHz, the Green Bank GB6 catalog (Gregory1996) at 4.85 GHz, the RACS333Rapid ASKAP Continuum Survey (Hale2021) at 888 MHz and the SUMSS444Sydney University/Molonglo Sky Survey. (Mauch2013) at 843 MHz.
@@
Because radio emission is not scattered or absorbed by dust grains and Compton thick structures are transparent to radio frequencies, this range yields complementary information to optical observations, tracing free-free and synchrotron processes, and the Rayleigh-Jeans tail of thermal emission, which are not accessible at other wavelengths (Condon2016; Padovani2016). Thus, radio surveys are indispensable for the study of a wide array of astronomical sources, including radio galaxies, quasars, ultra luminous starburst galaxies, Active Galactic Nuclei (AGN), planetary nebulae, supernova remnants, pulsars, and variable stars.
@@
Of particular interest for this work are low-frequency radio surveys (), which are sensitive to older synchrotron emission and therefore probe aged and extended radio structures where self-absorption effects are not significant, such as halos, relics, and jet lobes (Melrose1985; Ferrari2008). Some of the most notable low-frequency surveys are the VLA Low-frequency Sky Survey (VLSS) at 74 MHz (Cohen2007), the 8C radio survey at 38 MHz (Rees1990), the GaLactic and Extragalactic All-sky MWA survey (GLEAM) at 72-231 MHz (Wayth2015; Hurley2017) and the Westerbork Northern Sky Survey (WENSS) at 330 MHz (deBruyn2000). These cover a wide sky area with 0.5–4.5 arcmin resolution and flux density limits of 18–1000 mJy at . A breakthrough in sensitivity and sky coverage was achieved with the LOw Frequency ARray (LOFAR; Haarlem2013) Two-metre Sky Survey (LoTSS; Shimwell2022). LoTSS DR2 covers the 120–168 MHz range over 5 700 deg2 of the northern sky at 6 arcsec resolution, reaching a flux limit of 0.4 mJy for compact sources and a sensitivity of 0.25 mJy/beam at , providing an eightfold increase in the surface density of radio sources compared to previous works. This is complemented by the ongoing 42–66 MHz LOFAR Low Band Antenna Sky Survey (LoLSS; deGasperin2021) and the LOFAR Decametre Sky Survey (LoDSS) at 14–30 MHz. Some remarkable results are the discovery of a 7 Mpc giant radio galaxy (Oei24) or the search for stellar radio emission as a possible indicator of exoplanet interaction with M dwarfs (Callingham2021).
@@
The real potential of radio surveys is unleashed when they are combined with other frequencies that offer complementary information for a comprehensive understanding of astronomical sources. For instance, optical, infrared (IR), and ultraviolet observations enable the measurement of photometric and spectroscopic redshift, stellar masses, and star formation rates (SFRs) of galaxies. They also allow estimates the bolometric luminosity and radio-loudness of quasi-stellar objects (QSOs), as well as the mass and variability of stars. In particular, optical spectroscopy of radio sources provides more precise determination of the origin of the radio emission and the nature of the host. Examples of such multi-wavelength studies include the cross-match between SDSS555Sloan Digital Sky Survey (York2000; Stoughton2002; Strauss2002)., NVSS, and FIRST by Best2005 and the identification of NVSS sources within the 6dF Galaxy Survey (Mauch2007), which together show that radio-loud AGN preferentially reside in massive galaxies. In this regard, the co-evolution of SMBHs and galaxies has been investigated by Heckman2014, who combined radio, IR, optical, and X-ray data, and by Ching2017, who analyzed the LARGESS666Large Area Radio Galaxy Evolution Spectroscopic Survey. sample of FIRST sources with SDSS spectroscopy. These works revealed that the majority of radio AGN are Low-Excitation Radio Galaxies (LERGs), associated with jet activity from massive black holes in elliptical galaxies. By contrast, High-Excitation Radio Galaxies (HERGs) are associated with less massive, radiatively efficient black holes typically found in bluer, star-forming galaxies. In a follow-up study with the GAMA777Galaxy And Mass Assembly. survey, Ching2017b showed that luminous LERGs are preferentially associated with denser environments and galaxy groups. More recently, Arnaudova2024 combined LoTSS with spectroscopically confirmed QSOs from SDSS DR14 (Abolfathi2018) to show that radio-loud QSOs tend to exhibit higher SFRs than their radio-quiet counterparts.
@@
On the other hand, photometric surveys allows efficient access to extensive, large-area samples without the need for targeted observations. This enables photometric redshift (photo-) estimates, albeit with lower precision and potential contamination by outliers. To amend this, the use of narrow-band optical observations provides, among other, improved photo-z estimations by exploiting spectral features in galaxies, essential for clustering studies (Molino2019), and precise stellar parameter determinations (Yuan2023). Additionally, narrow-band filters offer detailed information on the spectral energy distribution (SED) and the emission lines for a better characterization of the star formation main sequence and SFR density in galaxies (VilellaRojo2021), the discovery of high-z Lyman QSOs (Spinoso2020), and the identification of large samples of extreme emission-line galaxies (Lumbreras2022). The catalog presented in this study exploits the advantages of narrow-band photometry from the Javalambre Photometric Local Universe Survey (J-PLUS DR3; Cenarro19; LopezSanJuan24) with low-frequency radio counterparts from the LOFAR Two-Metre Sky Survey (LoTSS DR2), and IR photometry from WISE. J-PLUS filters are specifically designed to provide accurate stellar population parameters and low-redshift photo- estimates. While other broad-band surveys such as DESI Legacy Imaging Surveys (Dey2019) offer deeper and wider coverage, J-PLUS adds the ability to resolve spectral features and provides improvements in the photo- and SFR estimations of high signal-to-noise sources, particularly when prominent spectral features such as bright emission lines or the Å break are present.
@@
The resulting product is the J-HERTz catalog (J-PLUS Heritage Exploration of Radio Targets at z 5), a public, multi-wavelength resource covering 2100 deg2 to characterize nearly half a million sources in the northern sky. J-HERTz provides a uniform, cross-matched data set with probabilistic Bayesian-neural-network classifications for galaxies, QSOs, and stars; improved photo- estimates that leverage the J-PLUS narrow bands; and a radio-loudness indicator based on LOFAR/WISE flux ratios, together with SED-based stellar masses and SFR for the galaxy population. These elements enable population-wide studies spanning galaxy evolution, AGN duty cycles and feedback, and the demographics of radio sources across astrophysical classes – including a large sample of candidate radio-emitting stars – within a single, coherent framework. These capabilities will significantly improve with the upcoming J-PAS survey (Bonoli2021, Vázquez-Ramió, in prep.), including 56 narrow bands and deeper observations with respect to J-PLUS by mag., which will be incorporated in future releases of J-HERTz. Throughout this paper, we adopt a cosmology with km s-1 Mpc-1, , and
2 A radio-optical catalog for the Northern Hemisphere
@@
2.1 Catalog description
The J-HERTz catalog contains objects found in J-PLUS DR3 and LoTSS DR2. The available information for each object includes IDs, optical fluxes and magnitudes, radio fluxes and sizes, WISE band magnitudes, photometric and spectroscopic redshifts and source classification. The catalog is accessible through the J-PLUS DR3 Data Access Services888https://archive.cefca.es/catalogues/jplus-dr3 and can be queried through TAP or ADQL services. All the included parameters are listed in the Appendix. We show the distribution of some of the main properties in Section 3 and discuss them in Section 4.
@@ 2.2 J-PLUS DR3
The J-PLUS DR3 catalog covers about of the northern sky, containing 47 million objects between galaxies, QSOs and stars. The observations are done at the Observatorio Astrofísico de Javalambre (Cenarro14) with the Javalambre Auxilary Survey Telescope of 83 cm diameter (JAST80) and T80Cam, a 9.2k 9.2k pixels CCD that provides a field of view and a pixel scale of arcsec pix-1 (MarinFranch15). It has 12 filters described in Table 1, a combination of 5 broad and 7 medium or narrow bands, reaches a depth of AB 21.8 mag at the r band at 5 with a 3 arcsec aperture and a arcsec full width at half maximum (FWHM) point spread function. J-PLUS provides flux densities and magnitudes in its 12 available bands, allowing for an accurate characterization of stars in our galaxy as well as 2D photo-spectral information for resolved galaxies in the local Universe (Rahna2025) and accurate photo-z estimates for moderately bright extra galactic sources. For our catalog and this work we use their AUTO photometry otherwise stated, which takes an automatically fitted ellipse from SExtractor. J-PLUS archives also provide cross-matches with other catalogs like Sloan Digital Sky Survey (SDSS) DR12 (Alam2015), catWISE2020 (Marocco2020) and Gaia DR3 (Gaia2023).
2.3 LoTSS DR2
@@
LoTSS DR2 (Shimwell2022) is a low-frequency radio catalog that provides fluxes at 144 MHz with a 6 arcsec resolution, which is particularly useful to detect wide area radio emission with lower surface brightness. It is based on LOFAR observations and covers 27% of the Northern Hemisphere sky ( 5700 deg2), containing around 4.4 million radio objects. It includes optical and/or IR identifications in unWISE (Lang2014) and DESI Legacy Imaging Surveys DR9 for 85% of the objects using a combination of likelihood-ratio methods, algorithmic optical identification and extensive visual inspection (Hardcastle2023). The catalog provides fitted magnitudes from WISE as well as other optical and UV catalogs, estimated radio emission sizes, photometric redshifts, and spectroscopic redshifts from SDSS DR16 (Ahumada2020), the first data release of the DESI spectroscopic survey (DESI2025) and the Hobby-Eberly Telescope Dark Energy Experiment (HETDEX) DR1 (Davis2023) when available. Using the most reliable redshifts, they also compute the linear sizes, radio luminosities, stellar masses and rest-frame magnitudes.
2.4 Radio-optical cross-match
@@
We use the optical/IR identifications of LoTSS provided by Hardcastle2023. In their work, they first cross-match the DESI Legacy Imaging Surveys to the unWISE catalog using a 2 arcsec radius to determine an optical/IR position, which is then assigned to a radio source in LoTSS DR2. We do not cross-match these identifications with the J-PLUS positions directly, but rather with the catWISE and SDSS identifications that J-PLUS provides. As most of the common objects in J-PLUS and LoTSS have identifications in WISE, this ensures that they are cross-matched using the same survey as reference and maximizes genuine matches. We first cross-match a subsample of the positions with a high tolerance (5 arcsec) and analyze the separations. The median is 0.23 arcsec, indicating a good agreement in most positions, while the 90th and 98th percentiles are 1.26 arcsec and 3.24 arcsec, respectively. We decide to use a 1 arcsec tolerance for the cross-match, which is slightly conservative to avoid spurious associations but still keeps a high completeness of genuine matches. We obtain 489,897 unique objects in total, which correspond to 80% of the LoTSS DR2 Hardcastle2023 identifications with a S/N 5 in the common area. The angular distances between the SDSS/catWISE identifications and the original J-PLUS positions are provided in the AngDist parameter, and 92% of them are smaller than 2 arcsec. Fig. 2 shows three examples of objects contained in the catalog, including J-PLUS 3 arcsec photometry, LoTSS radio emission and SDSS optical spectra. Panel (a) shows a star-forming galaxy, as can be seen from the strong H emission line. Panel (b) presents a QSO with C III] and C IV emission lines and a very strong blue continuum coming from the AGN. Finally, panel (c) is a radio galaxy whose jets extend 10 arcmin and has a drop in emission at blue wavelengths, indicating the presence of an old stellar population. There is an overall good agreement between the optical SED of J-PLUS and the optical spectra of SDSS of the examples, with small deviations that may be caused by different calibrations, PSF models and sky subtractions, or due to some lost flux outside of the SDSS fiber.
2.5 Object classification

### v2

@@ I Introduction
Radio astronomy has made substantial advances since the first radio emission coming from the center of our galaxy was discovered (Jansky, 1933). In 90 years, radio continuum surveys have improved from angular resolutions of several degrees to sub-arcsecond scales thanks to the development of radio interferometry, and from frequencies of a few hundred MHz to a full coverage of the radio window up to millimeter wavelengths. This advancement came hand in hand with the technical development that enabled the construction of larger and more efficient antennas, as well as improved backend systems with more sensitive receivers and faster correlators. Some of the most important radio continuum surveys in the last decades include the NVSS111NRAO (National Radio Astronomy Observatory) VLA Sky Survey. (Condon et al., 1998) at 1.4 GHz, the Cambridge 3C Radio Survey (Bennett, 1962) at 178 MHz, the Very Large Array (VLA) FIRST222Faint Images of the Radio Sky at Twenty-Centimeters. Survey (Becker et al., 1994) at 1.5 GHz, the Green Bank GB6 catalog (Gregory et al., 1996) at 4.85 GHz, the RACS333Rapid ASKAP Continuum Survey (Hale et al., 2021) at 888 MHz and the SUMSS444Sydney University/Molonglo Sky Survey. (Mauch et al., 2013) at 843 MHz.
@@
Because radio emission is not scattered or absorbed by dust grains and Compton thick structures are transparent to radio frequencies, this range yields complementary information to optical observations, tracing free-free and synchrotron processes, and the Rayleigh-Jeans tail of thermal emission, which are not accessible at other wavelengths (Condon and Ransom, 2016; Padovani, 2016). Thus, radio surveys are indispensable for the study of a wide array of astronomical sources, including radio galaxies, quasars, ultra luminous starburst galaxies, Active Galactic Nuclei (AGN), planetary nebulae, supernova remnants, pulsars, and variable stars.
@@
Of particular interest for this work are low-frequency radio surveys (), which are sensitive to older synchrotron emission and therefore probe aged and extended radio structures where self-absorption effects are not significant, such as halos, relics, and jet lobes (Melrose, 1985; Ferrari et al., 2008). Some of the most notable low-frequency surveys are the VLA Low-frequency Sky Survey (VLSS) at 74 MHz (Cohen et al., 2007), the 8C radio survey at 38 MHz (Rees, 1990), the GaLactic and Extragalactic All-sky MWA survey (GLEAM) at 72-231 MHz (Wayth et al., 2015; Hurley-Walker et al., 2017) and the Westerbork Northern Sky Survey (WENSS) at 330 MHz (de Bruyn et al., 2000). These cover a wide sky area with 0.5–4.5 arcmin resolution and flux density limits of 18–1000 mJy at . A breakthrough in sensitivity and sky coverage was achieved with the LOw Frequency ARray (LOFAR; van Haarlem et al., 2013) Two-metre Sky Survey (LoTSS; Shimwell et al., 2022). LoTSS DR2 covers the 120–168 MHz range over 5 700 deg2 of the northern sky at 6 arcsec resolution, reaching a flux limit of 0.4 mJy for compact sources and a sensitivity of 0.25 mJy/beam at , providing an eightfold increase in the surface density of radio sources compared to previous works. This is complemented by the ongoing 42–66 MHz LOFAR Low Band Antenna Sky Survey (LoLSS; de Gasperin et al., 2021) and the LOFAR Decametre Sky Survey (LoDSS) at 14–30 MHz. Some remarkable results are the discovery of a 7 Mpc giant radio galaxy (Oei et al., 2024) or the search for stellar radio emission as a possible indicator of exoplanet interaction with M dwarfs (Callingham et al., 2021).
@@
The real potential of radio surveys is unleashed when they are combined with other frequencies that offer complementary information for a comprehensive understanding of astronomical sources. For instance, optical, infrared (IR), and ultraviolet observations enable the measurement of photometric and spectroscopic redshift, stellar masses, and star formation rates (SFRs) of galaxies. They also allow estimates the bolometric luminosity and radio-loudness of quasi-stellar objects (QSOs), as well as the mass and variability of stars. In particular, optical spectroscopy of radio sources provides more precise determination of the origin of the radio emission and the nature of the host. Examples of such multi-wavelength studies include the cross-match between SDSS555Sloan Digital Sky Survey (York et al., 2000; Stoughton et al., 2002; Strauss et al., 2002)., NVSS, and FIRST by Best et al. (2005) and the identification of NVSS sources within the 6dF Galaxy Survey (Mauch and Sadler, 2007), which together show that radio-loud AGN preferentially reside in massive galaxies. In this regard, the co-evolution of SMBHs and galaxies has been investigated by Heckman and Best (2014), who combined radio, IR, optical, and X-ray data, and by Ching et al. (2017b), who analyzed the LARGESS666Large Area Radio Galaxy Evolution Spectroscopic Survey. sample of FIRST sources with SDSS spectroscopy. These works revealed that the majority of radio AGN are Low-Excitation Radio Galaxies (LERGs), associated with jet activity from massive black holes in elliptical galaxies. By contrast, High-Excitation Radio Galaxies (HERGs) are associated with less massive, radiatively efficient black holes typically found in bluer, star-forming galaxies. In a follow-up study with the GAMA777Galaxy And Mass Assembly. survey, Ching et al. (2017a) showed that luminous LERGs are preferentially associated with denser environments and galaxy groups. More recently, Arnaudova et al. (2024) combined LoTSS with spectroscopically confirmed QSOs from SDSS DR14 (Abolfathi et al., 2018) to show that radio-loud QSOs tend to exhibit higher SFRs than their radio-quiet counterparts.
@@
On the other hand, photometric surveys allows efficient access to extensive, large-area samples without the need for targeted observations. This enables photometric redshift (photo-) estimates, albeit with lower precision and potential contamination by outliers. To amend this, the use of narrow-band optical observations provides, among other, improved photo-z estimations by exploiting spectral features in galaxies, essential for clustering studies (Molino et al., 2019), and precise stellar parameter determinations (Yuan et al., 2023). Additionally, narrow-band filters offer detailed information on the spectral energy distribution (SED) and the emission lines for a better characterization of the star formation main sequence and SFR density in galaxies (Vilella-Rojo et al., 2021), the discovery of high-z Lyman QSOs (Spinoso et al., 2020), and the identification of large samples of extreme emission-line galaxies (Lumbreras-Calle et al., 2022). The catalog presented in this study exploits the advantages of narrow-band photometry from the Javalambre Photometric Local Universe Survey (J-PLUS DR3; Cenarro et al., 2019; López-Sanjuan et al., 2024) with low-frequency radio counterparts from the LOFAR Two-Metre Sky Survey (LoTSS DR2), and IR photometry from WISE. J-PLUS filters are specifically designed to provide accurate stellar population parameters and low-redshift photo- estimates. While other broad-band surveys such as DESI Legacy Imaging Surveys (Dey et al., 2019) offer deeper and wider coverage, J-PLUS adds the ability to resolve spectral features and provides improvements in the photo- and SFR estimations of high signal-to-noise sources, particularly when prominent spectral features such as bright emission lines or the Å break are present.
@@
The resulting product is the J-HERTz catalog (J-PLUS Heritage Exploration of Radio Targets at z 5), a public, multi-wavelength resource covering 2100 deg2 to characterize nearly half a million sources in the northern sky. J-HERTz provides a uniform, cross-matched data set with probabilistic Bayesian-neural-network classifications for galaxies, QSOs, and stars; improved photo- estimates that leverage the J-PLUS narrow bands; and a radio-loudness indicator based on LOFAR/WISE flux ratios, together with SED-based stellar masses and SFR for the galaxy population. These elements enable population-wide studies spanning galaxy evolution, AGN duty cycles and feedback, and the demographics of radio sources across astrophysical classes – including a large sample of candidate radio-emitting stars – within a single, coherent framework. These capabilities will significantly improve with the upcoming J-PAS survey (Bonoli et al., 2021, Vázquez-Ramió, in prep.), including 56 narrow bands and deeper observations with respect to J-PLUS by mag., which will be incorporated in future releases of J-HERTz. Throughout this paper, we adopt a cosmology with km s-1 Mpc-1, , and
II A radio-optical catalog for the Northern Hemisphere
@@
II.1 Catalog description
The J-HERTz catalog contains objects found in J-PLUS DR3 and LoTSS DR2. The available information for each object includes IDs, optical fluxes and magnitudes, radio fluxes and sizes, WISE band magnitudes, photometric and spectroscopic redshifts and source classification. The whole catalog is available in the online version of this publication and in Zenodo at doi:10.5281/zenodo.17651891. Additionally, a reduced version with the object IDs and some of the value-added parameters is accessible through the J-PLUS DR3 Data Access Services888https://archive.cefca.es/catalogues/jplus-dr3 and can be queried through TAP or ADQL services. All the included parameters are listed in the Appendix. We show the distribution of some of the main properties in Section III and discuss them in Section IV.
@@ II.2 J-PLUS DR3
The J-PLUS DR3 catalog covers about of the northern sky, containing 47 million objects between galaxies, QSOs and stars. The observations are done at the Observatorio Astrofísico de Javalambre (Cenarro et al., 2014) with the Javalambre Auxilary Survey Telescope of 83 cm diameter (JAST80) and T80Cam, a 9.2k 9.2k pixels CCD that provides a field of view and a pixel scale of arcsec pix-1 (Marín-Franch et al., 2015). It has 12 filters described in Table 1, a combination of 5 broad and 7 medium or narrow bands, reaches a depth of AB 21.8 mag at the r band at 5 with a 3 arcsec aperture and a arcsec full width at half maximum (FWHM) point spread function. J-PLUS provides flux densities and magnitudes in its 12 available bands, allowing for an accurate characterization of stars in our galaxy as well as 2D photo-spectral information for resolved galaxies in the local Universe (Rahna et al., 2025) and accurate photo-z estimates for moderately bright extra galactic sources. For our catalog and this work we use their AUTO photometry otherwise stated, which takes an automatically fitted ellipse from SExtractor. J-PLUS archives also provide cross-matches with other catalogs like Sloan Digital Sky Survey (SDSS) DR12 (Alam et al., 2015), catWISE2020 (Marocco et al., 2020) and Gaia DR3 (Gaia Collaboration et al., 2023).
II.3 LoTSS DR2
@@
LoTSS DR2 (Shimwell et al., 2022) is a low-frequency radio catalog that provides fluxes at 144 MHz with a 6 arcsec resolution, which is particularly useful to detect wide area radio emission with lower surface brightness. It is based on LOFAR observations and covers 27% of the Northern Hemisphere sky ( 5700 deg2), containing around 4.4 million radio objects. It includes optical and/or IR identifications in unWISE (Lang, 2014) and DESI Legacy Imaging Surveys DR9 for 85% of the objects using a combination of likelihood-ratio methods, algorithmic optical identification and extensive visual inspection (Hardcastle et al., 2023). The catalog provides fitted magnitudes from WISE as well as other optical and UV catalogs, estimated radio emission sizes, photometric redshifts, and spectroscopic redshifts from SDSS DR16 (Ahumada et al., 2020), the first data release of the DESI spectroscopic survey (DESI Collaboration et al., 2025) and the Hobby-Eberly Telescope Dark Energy Experiment (HETDEX) DR1 (Davis et al., 2023) when available. Using the most reliable redshifts, they also compute the linear sizes, radio luminosities, stellar masses and rest-frame magnitudes.
II.4 Radio-optical cross-match
@@
We use the optical/IR identifications of LoTSS provided by Hardcastle et al. (2023). In their work, they first cross-match the DESI Legacy Imaging Surveys to the unWISE catalog using a 2 arcsec radius to determine an optical/IR position, which is then assigned to a radio source in LoTSS DR2. We do not cross-match these identifications with the J-PLUS positions directly, but rather with the catWISE and SDSS identifications that J-PLUS provides. As most of the common objects in J-PLUS and LoTSS have identifications in WISE, this ensures that they are cross-matched using the same survey as reference and maximizes genuine matches. We first cross-match a subsample of the positions with a high tolerance (5 arcsec) and analyze the separations. The median is 0.23 arcsec, indicating a good agreement in most positions, while the 90th and 98th percentiles are 1.26 arcsec and 3.24 arcsec, respectively. We decide to use a 1 arcsec tolerance for the cross-match, which is slightly conservative to avoid spurious associations but still keeps a high completeness of genuine matches. We obtain 489,897 unique objects in total, which correspond to 80% of the LoTSS DR2 Hardcastle et al. (2023) identifications with a S/N 5 in the common area. The angular distances between the SDSS/catWISE identifications and the original J-PLUS positions are provided in the AngDist parameter, and 92% of them are smaller than 2 arcsec. Fig. 2 shows three examples of objects contained in the catalog, including J-PLUS 3 arcsec photometry, LoTSS radio emission and SDSS optical spectra. Panel (a) shows a star-forming galaxy, as can be seen from the strong H emission line. Panel (b) presents a QSO with C III] and C IV emission lines and a very strong blue continuum coming from the AGN. Finally, panel (c) is a radio galaxy whose jets extend 10 arcmin and has a drop in emission at blue wavelengths, indicating the presence of an old stellar population. There is an overall good agreement between the optical SED of J-PLUS and the optical spectra of SDSS of the examples, with small deviations that may be caused by different calibrations, PSF models and sky subtractions, or due to some lost flux outside of the SDSS fiber.
II.5 Object classification